# Yunjue Agent Tech Report: A Fully Reproducible, Zero-Start In-Situ Self-Evolving Agent System for Open-Ended Tasks

### v1

@@ 5.2 Warm start evolution in shifting domains
Figure 4 illustrates the evolution of the tool library size relative to the cumulative number of processed queries. The overall trajectory demonstrates a global trend toward convergence. A localized surge in tool synthesis during the late phase of HLE (queries 2,000–2,400) is driven by intra-benchmark domain shifts: as query semantics transition from predominantly mathematical to social sciences and other disciplines, a targeted expansion of the toolset is necessitated. Notably, despite this semantic heterogeneity, only 97 tools were generated across the entire 2,500-HLE query corpus, evidencing the effective consolidation of acquired knowledge. Subsequently, the tool growth curves for DSQA and xSciQA exhibit a near-zero gradient. This plateau indicates that the capabilities evolved during the HLE phase possess sufficient generalization to support robust adaptation to novel domains with negligible marginal tool synthesis.
To empirically substantiate the role of tools as vehicles for knowledge crystallization and domain adaptation, we analyze the toolset intersection visualized in Figure 5. The comparative analysis reveals a substantial overlap between tools generated in the zero-start regime and those synthesized during warm-start evolution. Specifically, the recurrence of exact matches, such as read_excel, validates the system’s capacity to deterministically recover essential utilities. Furthermore, the emergence of semantically aligned variants (e.g., inspect_ods, fetch_api) highlights the agent’s adaptability in synthesizing functionally equivalent primitives tailored to specific task constraints. The predominance of these convergent tools corroborates our hypothesis of systemic robustness in shifting domains, confirming that the agent reliably gravitates toward an invariant core of capabilities to fulfill domain requirements, independent of the initialization state.
@@ 5.3 Evolutionary generality loss
To quantify the dynamics of knowledge acquisition, we propose a novel metric termed Evolutionary generality loss (EGL). Functioning analogously to the objective function in stochastic optimization, EGL serves as a real-time indicator of evolutionary convergence. Formally, EGL is defined as:
$\text{EGL}=\frac{\sum_{i=1}^{N}c_{i}}{\sum_{i=1}^{N}u_{i}}\times 1000$ (1)
where $c_{i}$ and $u_{i}$ denote the number of newly synthesized tools and tool invocations for the $i$-th query, respectively, and $N$ represents the total number of processed queries. Intuitively, the EGL value is elevated during the nascent stages of evolution, driven by the necessity of ab initio tool creation. As the system matures, the dominance of tool reuse over synthesis causes the metric to decay. Conversely, a sustained high EGL value implies that the system is failing to generalize, necessitating continuous creation. Conceptually, this metric is inversely correlated with tool generality: as the versatility of the repository increases, the EGL value decreases. To validate its efficacy, we monitored the EGL trajectory on the HLE benchmark as a function of query volume. The resulting trend in Figure 6 exhibits a distinct convergence pattern, achieving asymptotic stability after processing approximately 1,000 queries (40% of the stream).
Figure 7: Evolution dynamics under different batch sizes. (a) The number of created tools with the number of queries. (b) Average token consumption per tool use. (c) Cumulative number of tool invocations.
To empirically corroborate this convergence, we extracted system snapshots at 10%, 40%, 70%, and 100% of the evolutionary timeline and evaluated performance on a stratified random sample of 200 queries (100 each from the HLE and DSQA datasets), with further tool synthesis disabled. As illustrated in the bar chart of Figure 6, a substantial performance inflection occurs between the 10% and 40% checkpoints, signifying a rapid expansion of the system’s capability boundary. Conversely, marginal gains diminish significantly in the 40%–100% interval. This saturation provides strong evidence that the evolutionary process successfully converges, reaching a state of high generality where incremental tool synthesis yields diminishing returns.
Table 2: Experimental results on a subset of 200 selected queries under varying batch sizes. “Tool succ. rt.” denotes the tool execution success rate, while “Avg #tokens for tools” quantifies the average token consumption per tool invocation. The condition $\infty$” represents that the backend model operates exclusively with a Python interpreter, processing queries independently without tool accumulation.
Batch size Accuracy Tool succ. rt. Avg #tokens for tools
2 48.0 99.6% 109.9
4 50.0 99.9% 130.3
8 47.5 99.6% 188.4
16 51.5 99.1% 108.4
$\infty$ 40.0 81.8% 518.2
@@ 5.4 Ablation study
To elucidate the factors underpinning the efficacy of tool evolution and to ablate key system parameters, we conduct comparative experiments using the identical subset of 200 queries employed in the EGL evaluation.
Why does tool evolution work? To validate the fundamental utility of tool evolution, we establish a baseline comparison against an agent sharing an identical backend (GPT-5) but equipped solely with a Python interpreter. While the interpreter allows for ad-hoc code execution and iterative debugging based on error feedback, it fundamentally lacks the capacity for knowledge crystallization and reuse.
We evaluate our agent under a zero-start regime with varying batch sizes, contrasting it against the Python interpreter baseline (denoted as $B\to\infty$). The condition $\infty$” represents the theoretical asymptotic limit of infinite batch size, where the backend model operates exclusively with a Python interpreter. In this setting, queries are processed independently, precluding any tool accumulation, clustering, or shared knowledge transfer. As detailed in Table 2, the Python-only baseline yields a markedly inferior score of 40.0, compared to the 47.5–51.5 range achieved by our method. To elucidate the underlying causes, we analyzed the tool success rate (successful invocations per total attempts) and token efficiency (average tokens per invocation). The baseline demonstrates deficiencies in both metrics, exhibiting a success rate of only 81.8% while consuming an average of 518.2 tokens per invocation. In stark contrast, our system consistently maintains success rates exceeding 99% with significantly reduced token overhead. This divergence underscores that evolved tools are not only more robust but also more efficient in information processing. We posit that the high execution error rate and the verbosity of raw code generation induce severe contextual contamination, thereby degrading the reasoning capabilities of the Python baseline.
Table 3: Performance comparison and number of created tools of different backend models on DeepSearchQA and FinSearchComp benchmarks.
@@
How does batch size impact parallel evolution? We further investigate the role of batch size $B$ in mediating the trade-off between evolutionary velocity and convergence stability. As illustrated in Figure 7(a), larger batch configurations induce a steeper initial gradient in tool synthesis. This suggests that high-concurrency processing facilitates the rapid identification of common patterns and the accelerated crystallization of foundational primitives. Crucially, despite varied initial velocities, the terminal library sizes remain comparable across all settings, confirming that the system invariably converges to a consistent equilibrium of capabilities regardless of batch magnitude.
Figure 7(b) tracks the temporal evolution of computational overhead. Even in the nascent stages, our parallelized architecture yields reduced token consumption relative to the baseline (w/ py exec). As the system matures, the average token cost per invocation demonstrates a precipitous decay, eventually stabilizing at an asymptotic minimum. Note that localized fluctuations in this downward trend correspond to the sporadic synthesis of novel tools required for outlier queries. This overall convergence signifies the system’s successful phase transition from expensive ab initio code generation to efficient, low-latency tool reuse.
We observe in Figure 7(b) that the average token consumption per tool use after convergence for a batch size of 8 is slightly higher than other settings. This is because, as shown in Figure 7(a), it generates the largest number of tools, and Figure 7(c) indicates it has the highest cumulative tool invocations. We attribute this to the stochastic nature of LLM execution, which leads to varying agent behaviors under different settings. Future work will investigate strategies to align the convergence states of the system across different batch sizes.
How does backend model selection affect evolution? We evaluate the influence of the underlying foundation model on system performance across the HLE and DSQA datasets, as detailed in Table 3. Results indicate that when instantiated with GPT-5, our agent surpasses the majority of established baselines. Notably, the lightweight GPT-5-mini variant achieves competitive performance, demonstrating significant utility despite its reduced parameter count. Crucially, we employ a unified prompt strategy across all configurations without architecture-specific tuning. These findings underscore the model agnosticism and generalization capability of our framework, confirming its robustness across diverse backend scales.
@@ 6 Related Work
Self-evolving agents. To transcend the limitations of static designs, general-purpose agent systems must adapt to open-ended, interactive environments in real-time (Gao et al., 2025). Self-evolving agents achieve this by autonomously refining internal components, categorized into workflow optimization, model parameter updates, context management, and tool synthesis. Specifically, MAS-GPT (Ye et al., 2025) and AFlow (Zhang et al., 2025a) generate query-specific workflows, while TT-SI (Acikgoz et al., 2025) and SCA (Zhou et al., 2025) update parameters by generating challenging training problems. Furthermore, ELL (Cai et al., 2025) and Zhang et al. (2025b) evolve by accumulating environmental interaction experience as context. Regarding tool synthesis, LIVE-SWE-agent (Xia et al., 2025) and STELLA (Jin et al., 2025) enable on-the-fly tool creation for software engineering and biomedical research, respectively. Nevertheless, most self-evolving methods necessitate explicit training. Crucially, even approaches closest to our work—those autonomously creating tools at test-time—remain confined to specific domains and lack tool reuse mechanisms, hindering their general applicability.
@@
Tool evolution agents. The paradigm of autonomous tool synthesis has garnered significant attention in recent literature. Pioneering systems such as Voyager (Wang et al., 2023), STELLA (Jin et al., 2025), and LIVE-SWE-agent (Xia et al., 2025) empower agents to generate executable tools for embodied control or software engineering tasks. Notably, analyses of the tool embedding space in LIVE-SWE-agent (Xia et al., 2025) reveal distinct clustering patterns among functional equivalents; this observation directly informed the design of our absorbing mechanism, which utilizes clustering to consolidate redundant utilities. Regarding tool reuse, Alita (Qiu et al., 2025) facilitates the persistence of successful tools, though its optimization trajectory is predicated on ground-truth supervision. In the context of inference-time adaptation, Lu et al. (2026) proposed a test-time tool evolution framework governed by a cost-sensitive objective function, yet its application remains narrowly scoped to scientific reasoning. Collectively, these approaches prioritize the optimization of existing assets or domain-specific synthesis, distinguishing them from our framework’s focus on self-evolution in open-ended environments.
@@
General agent systems. Recent research has sought to construct general-purpose agents capable of cross-domain operation. DeepAgent (Li et al., 2025) utilizes a scalable toolset to address diverse tasks but lacks the capacity for autonomous environmental adaptation, leading to performance limitations. Agent KB (Tang et al., 2025) attempts to solve cross-domain problems by summarizing execution experiences from other agents; however, its performance is constrained by a deficiency in necessary tools. In contrast, our approach achieves generality by continuously creating, reusing, and optimizing tools in-situ, effectively overcoming the limitations of static toolsets and experience-based adaptation.
@@ 8 Conclusion
In this work, we presented the In-Situ Self-Evolving framework, enabling LLM-based agents to autonomously adapt and evolve within open-ended environments. Through the Yunjue Agent and the proposed Parallel Batch Evolution strategy, we demonstrated that treating tools as dynamic vehicles for knowledge crystallization allows for robust zero-start learning and the emergence of generalized, transferable capabilities. Our empirical results validate that this approach not only achieves state-of-the-art performance across heterogeneous benchmarks but also significantly outperforms static baselines in both efficiency and adaptability. Furthermore, warm-start experiments confirm that the accumulated generalized knowledge can be seamlessly transferred to novel domains, enabling the agent to adapt to shifting environments with minimal marginal tool synthesis. We release our codebase, evaluation and system traces to the open-source community to facilitate further research. Broadly, future work will aim to unify tool evolution with memory and workflow adaptation, advancing toward more autonomous and general-purpose agentic systems.
Appendix A Experiment setup

### v2

@@ 5.2 Warm-start evolution in shifting domains
Figure 4: Venn diagram comparing unique zero-start tools ($\mathcal{T}_{\text{DSQA}}\setminus\mathcal{T}_{\text{HLE}}$, left) and incremental warm-start tools ($\mathcal{T}_{\text{HLE}\rightarrow\text{DSQA}}\setminus\mathcal{T}_{\text{HLE}}$, right) on DeepSearchQA. Points represent semantically clustered tools. The intersection (71%) denotes functional overlap, with dashed lines marking exact matches, while the difference set accounts for 29%.
Figure 5 illustrates the evolution of the tool library size relative to the cumulative number of processed queries. The overall trajectory demonstrates a global trend toward convergence. A localized surge in tool synthesis during the late phase of HLE (queries 2,000–2,400) is driven by intra-benchmark domain shifts: as query semantics transition from predominantly mathematical to social sciences and other disciplines, a targeted expansion of the toolset is necessitated. Notably, despite this semantic heterogeneity, only 97 tools were generated across the entire 2,500-HLE query corpus, evidencing the effective consolidation of acquired knowledge. Subsequently, tool growth trajectories across secondary datasets exhibit near-zero slopes, indicating that HLE-evolved capabilities facilitate robust adaptation to novel domains with negligible marginal synthesis.
To substantiate tools as vehicles for knowledge crystallization and domain adaptation, Figure 4 visualizes the intersection between zero-start and warm-start toolsets. The recurrence of exact matches (e.g., read_excel) validates the capacity to deterministically recover essential utilities. Furthermore, semantically aligned variants (e.g., seed_http_request vs. fetch_api_data, see Appendix E) highlight adaptive functional equivalence. This convergence confirms the agent gravitates toward an invariant core of capabilities independent of initialization.
Figure 6: EGL dynamics on HLE and accuracy on selected datasets vs. evolution step. The orange curve shows the EGL trend (left axis, scaled by 1000). Blue bars indicate the accuracy (right axis) of agents using HLE toolsets frozen at 10%, 40%, 70%, and 100% of evolution.
@@ 5.3 Evolutionary generality loss
To quantify knowledge acquisition, we introduce Evolutionary Generality Loss (EGL), a real-time convergence indicator analogous to training loss. EGL is defined as the cumulative ratio of synthesized tools to invocations:
$\text{EGL}(t)\triangleq\mathcal{L}_{t}=\frac{C_{t}}{U_{t}}=\frac{\sum_{i=1}^{t}c_{i}}{\sum_{i=1}^{t}u_{i}}$ (1)
where $c_{i}$ and $u_{i}$ denote tool synthesis and invocations at step $i$. EGL is inversely correlated with repository generality. As shown in Figure 6 (HLE dataset), initial high values reflect rapid accumulation, which stabilizes asymptotically after $\sim$1,000 queries (40% of the stream).
Figure 7: Evolutionary dynamics across varying batch sizes. (a) Cumulative tool synthesis relative. (b) Mean token consumption per invocation. (c) Aggregate count of tool invocations.
Appendix A provides the complete formal proofs. We first derive the sufficient condition for the monotonic decrease of EGL:
Theorem 5.1 (Descent Condition).
If the marginal synthesis ratio is strictly lower than the cumulative ratio (i.e., $\frac{c_{t+1}}{u_{t+1}}<\frac{C_{t}}{U_{t}}$), the EGL metric decreases ($\mathcal{L}_{t+1}<\mathcal{L}_{t}$).
Thus, EGL declines whenever the marginal creation rate falls below the historical average. Empirically (Figure 6), the trajectory is consistently downward, with transient spikes reflecting domain-specific bursts.
Second, the convergence rate:
Theorem 5.2 (Convergence Rate).
If cumulative tools stabilize ($C_{t}\to C_{\text{max}}$) and invocations grow linearly ($U_{t}\propto t$), EGL converges to zero ($\mathcal{L}_{t}\to 0$) at a rate of $\mathcal{O}(1/t)$.
Figures 5 and 7(c) empirically validate the stabilization of $C_{t}$ and linear growth of $U_{t}$, respectively.
To further validate convergence, we evaluated toolset snapshots (frozen at 10%, 40%, 70%, 100% of the timeline) on 200 stratified queries with further tool synthesis disabled. Figure 6 bar chart reveals a sharp performance inflection between 10%–40%, followed by diminishing returns from 40% to 100%. This saturation confirms the system successfully converges to a high-generality state.
Table 2: Ablation results on 200 queries across varying batch sizes (BS). Metrics include Accuracy (Acc, %), total inference time in minutes (Time), Tool Success Rate (TSR, %), Average Tokens per Tool (ATT), and cost per query (Cost, \$). The condition $\infty$” denotes a Python-interpreter baseline processing queries independently without tool accumulation.
BS Acc Time TSR ATT Cost
[…]
@@ 5.4 Ablation study
To investigate the efficacy of tool evolution, we conduct ablation studies using the 200-query EGL subset.
Mechanisms of tool evolution efficacy. We benchmark our approach against a Python-interpreter with the same baseline (GPT-5), representing the asymptotic limit of an infinite batch size ($B\to\infty$) where queries are processed independently without tool accumulation. As detailed in Table 2, this baseline demonstrates markedly inferior performance (40.0) compared to our method (47.5–51.5). Analysis suggests this performance gap stems from the baseline’s lower execution success rate (81.8% vs. >>99%) and significantly higher token consumption (518 tokens per invocation). We posit that the resulting verbosity and frequent execution failures induce contextual contamination, which subsequently degrades reasoning capabilities (see Appendix E).
Table 3: Performance comparison and tool library size of different backend models on DeepSearchQA and FinSearchComp.
@@
Impact of batch size on parallel evolution. We investigate the influence of batch size $B$ on evolutionary dynamics, evaluating configurations where $B\in\{2,4,8,16\}$ against a Python interpreter baseline ($B\to\infty$). As detailed in Table 2, all batch-evolved agents surpass the baseline, reinforcing the efficacy of tool accumulation. Furthermore, the results indicate that increasing $B$ proportionally reduces processing latency and marginally enhances accuracy without incurring substantial overhead, thereby validating the Parallel Batch Evolution strategy.
The impact of batch size on evolutionary trajectories is further analyzed in Figure 7(a). While larger batches accelerate initial tool crystallization, terminal library sizes remain comparable, suggesting the system converges toward a consistent capability equilibrium independent of $B$. Regarding computational efficiency, Figure 7(b) demonstrates that token expenditures consistently remain below the baseline, decaying toward an asymptotic minimum as the system transitions from costly ab initio tool generation to efficient reuse. A minor anomaly observed for $B=8$ in Table 2—characterized by higher token consumption—is attributed to maximal tool generation and invocation counts (Figure 7(a,c)) resulting from LLM stochasticity. Future research will explore methods to align convergence states across varying batch sizes.
Influence of backend model selection. We evaluate the impact of the underlying foundation model on system performance across the DSQA and FinSearchComp datasets (see Table 3). The results indicate that when instantiated with GPT-5, our agent surpasses the majority of established baselines. Notably, the lightweight GPT-5-mini variant achieves competitive performance, demonstrating significant utility despite its reduced parameter scale. These findings underscore the model-agnostic nature and generalization capabilities of the proposed framework, confirming its robustness across diverse backend architectures and scales.
@@ 6 Related work
Self-evolving agents. To transcend the limitations of static designs, general-purpose agent systems must adapt to open-ended, interactive environments in real-time (Gao et al., 2025). Self-evolving agents achieve this by autonomously refining internal components, categorized into workflow optimization, model parameter updates, context management, and tool synthesis. Specifically, MAS-GPT (Ye et al., 2025) and AFlow (Zhang et al., 2025b) generate query-specific workflows, while TT-SI (Acikgoz et al., 2025) and SCA (Zhou et al., 2025) update parameters by generating challenging training problems. Furthermore, ELL (Cai et al., 2025) and Zhang et al. (2025c) evolve by accumulating environmental interaction experience as context. Regarding tool synthesis, LIVE-SWE-agent (Xia et al., 2025) and STELLA (Jin et al., 2025) enable on-the-fly tool creation for software engineering and biomedical research, respectively. Nevertheless, most self-evolving methods necessitate explicit training. Crucially, even approaches closest to our work—those autonomously creating tools at test-time—remain confined to specific domains and lack tool reuse mechanisms, hindering their general applicability.
@@
Tool evolution agents. The paradigm of autonomous tool synthesis has garnered significant attention in recent literature. Pioneering systems such as Voyager (Wang et al., 2023), STELLA (Jin et al., 2025), and LIVE-SWE-agent (Xia et al., 2025) empower agents to generate executable tools for embodied control or software engineering tasks. Notably, analyses of the tool embedding space in LIVE-SWE-agent (Xia et al., 2025) reveal distinct clustering patterns among functional equivalents; this observation directly informed the design of our absorbing mechanism, which utilizes clustering to consolidate redundant utilities. Regarding tool reuse, Alita-G (Qiu et al., 2025) facilitates the persistence of successful tools, though its optimization trajectory is predicated on ground-truth supervision. In the context of inference-time adaptation, Lu et al. (2026) proposed a test-time tool evolution framework governed by a cost-sensitive objective function, yet its application remains narrowly scoped to scientific reasoning. Collectively, these approaches prioritize the optimization of existing assets or domain-specific synthesis, distinguishing them from our framework’s focus on self-evolution in open-ended environments.
@@
Agent systems for open-ended tasks. Recent research has sought to construct general-purpose agents capable of cross-domain operation. DeepAgent (Li et al., 2025) utilizes a scalable toolset to address diverse tasks but lacks the capacity for autonomous environmental adaptation, leading to performance limitations. Agent KB (Tang et al., 2025) attempts to solve cross-domain problems by summarizing execution experiences from other agents; however, its performance is constrained by a deficiency in necessary tools. In contrast, our approach achieves generality by continuously creating, reusing, and optimizing tools in-situ, effectively overcoming the limitations of static toolsets and experience-based adaptation.
@@ 8 Conclusion
In this work, we presented the In-Situ Self-Evolving framework, enabling LLM-based agents to autonomously adapt and evolve within open-ended environments. Through the Yunjue Agent and the proposed Parallel Batch Evolution strategy, we demonstrated that treating tools as dynamic vehicles for knowledge crystallization allows for robust zero-start learning and the emergence of generalized, transferable capabilities. Our empirical results validate that this approach achieves state-of-the-art performance across heterogeneous benchmarks. Furthermore, warm-start experiments confirm that the accumulated generalized knowledge can be seamlessly transferred to novel domains, enabling the agent to adapt to shifting environments. We release our codebase, evaluation and system traces to the open-source community to facilitate further research. Broadly, future work will aim to unify tool evolution with memory and workflow adaptation, advancing toward more autonomous and general-purpose agentic systems.
Appendix A Mathematical proof
This section details the formal proofs for Theorems 5.1 and 5.2, originally presented in Section 5.3.
A.1 Proof of theorem 5.1
We begin by establishing the sufficient condition for the monotonic decrease of the EGL metric at step $t+1$ (i.e., $\mathcal{L}_{t+1}<\mathcal{L}_{t}$). By definition:
$\mathcal{L}_{t+1}=\frac{C_{t}+c_{t+1}}{U_{t}+u_{t+1}}<\frac{C_{t}}{U_{t}}=\mathcal{L}_{t}$ (2)
Since the total invocation count $U_{t}$ is strictly positive, we cross-multiply without reversing the inequality:
$(C_{t}+c_{t+1})\cdot U_{t}<C_{t}\cdot(U_{t}+u_{t+1})$ (3)
Expanding the terms yields:
$C_{t}U_{t}+c_{t+1}U_{t}<C_{t}U_{t}+C_{t}u_{t+1}$ (4)
Subtracting the common term $C_{t}U_{t}$ and rearranging results in the Descent Condition:
$\frac{c_{t+1}}{u_{t+1}}<\frac{C_{t}}{U_{t}}$ (5)
This derivation demonstrates that $\mathcal{L}_{t}$ decreases if and only if the marginal ratio of synthesis to invocation at the current step ($c_{t+1}/u_{t+1}$) is strictly lower than the historical cumulative ratio ($C_{t}/U_{t}$). During early evolutionary stages, extensive exploration ($c_{t+1}>0$) may induce fluctuations. However, as the agent transitions to an exploitation phase ($c_{t+1}\to 0$ while $u_{t+1}\geq 1$), the marginal ratio vanishes, guaranteeing satisfaction of the inequality and ensuring the strict decrease of the metric.
A.2 Proof of theorem 5.2
To establish the asymptotic dynamics of $\mathcal{L}_{t}$, we introduce the following assumptions governing the evolutionary process:
Assumption A.1 (Knowledge Boundedness).
The task domain admits a finite set of optimal primitives sufficient to resolve all potential queries. As the agent evolves, the cumulative count of synthesized tools $C_{t}$ converges to a finite upper bound $C_{\text{max}}$. Consequently, marginal tool generation vanishes asymptotically:
$\lim_{t\to\infty}C_{t}=C_{\text{max}}\implies\lim_{t\to\infty}c_{t}=0$ (6)
We posit this as a mild condition. Empirically, as validated in Figure 6, the marginal growth of synthesized tools diminishes toward zero once the agent consolidates sufficient generalized knowledge.
Assumption A.2 (Continuous Activity).
We assume every newly synthesized tool is immediately invoked at least once, ensuring $0\leq c_{t}\leq u_{t}$ and $0\leq\frac{c_{t}}{u_{t}}\leq 1$. Operationally, if a synthesized tool remains uninvoked, it is discarded post-processing to satisfy this constraint. Additionally, we assume every query necessitates at least one tool invocation ($u_{t}\geq 1$), with the invocation count fluctuating around a stationary mean $\mu\geq 1$. This implies that cumulative invocations $U_{t}$ grow linearly with respect to $t$:
$U_{t}\approx\mu\cdot t\quad\text{as }t\to\infty$ (7)
Note that for the boundary condition $u_{t}=0$, we define $c_{t}/u_{t}=0$ to avoid singularity. Figure 7(c) provides empirical evidence supporting the linear growth of $U_{t}$ throughout the evolutionary steps.
Based on Assumptions A.1 and A.2, we analyze the limit of $\mathcal{L}_{t}$ as $t\to\infty$:
$\lim_{t\to\infty}\mathcal{L}_{t}=\frac{\lim_{t\to\infty}C_{t}}{\lim_{t\to\infty}U_{t}}\approx\lim_{t\to\infty}\frac{C_{\text{max}}}{\mu\cdot t}$ (8)
Since $C_{\text{max}}$ and $\mu$ are positive constants:
$\lim_{t\to\infty}\mathcal{L}_{t}=0$ (9)
This confirms that EGL functions mathematically as a loss metric, converging to zero at a rate of $\mathcal{O}(1/t)$ as system stability is achieved.
Appendix B Algorithm
Algorithm 1 Yunjue Agent Execution and Evolution Pipeline
1:Input: Query Batch $\mathcal{Q}_{t}=\{q_{t,1},\dots,q_{t,B}\}$, Global Tool Library $\mathcal{T}_{t-1}$
2:Output: Answers $\mathcal{A}_{t}$, Updated Library $\mathcal{T}_{t}$
[…]
4:for each query $q_{t,i}\in\mathcal{Q}_{t}$ in parallel do
5:  $\mathcal{R}_{miss}\leftarrow\emptyset$, $\text{done}\leftarrow\textbf{false}$
6:  while not done do
7:   $\omega_{req}\leftarrow\textbf{Manager}(q_{t,i},\mathcal{T}_{t-1},\mathcal{R}_{miss})$ {Analysis required tools}
8:   if new tool required then
9:    $\mathcal{P}_{t,i}\leftarrow\textbf{ToolDeveloper}(q_{t,i},\omega_{req})$ {Synthesize new tools}
10:    $\mathcal{T}_{local}\leftarrow\mathcal{T}_{t-1}\cup\mathcal{P}_{t,i}$
11:    $\mathcal{T}_{new}\leftarrow\mathcal{T}_{new}\cup\mathcal{P}_{t,i}$
12:   else
13:    $\mathcal{T}_{local}\leftarrow\mathcal{T}_{t-1}$
14:   end if
15:   $r_{t,i},\mathcal{R}_{miss}\leftarrow\textbf{Executor}(q_{t,i},\mathcal{T}_{local})$ {Execute and update capability missing report}
16:   if $\mathcal{R}_{miss}=\emptyset$ then
17:    $\text{done}\leftarrow\textbf{true}$
18:   end if
19:  end while
20:  $a_{t,i}\leftarrow\textbf{Integrator}(q_{t,i},r_{t,i})$ {Integrate final answer}
21:  $\mathcal{A}_{t}\leftarrow\mathcal{A}_{t}\cup\{a_{t,i}\}$
22:end for
[…]
25:return $\mathcal{A}_{t}$, $\mathcal{T}_{t}$
Appendix C Experiment setup